OpenStudy (anonymous):

If 6x - 3y = 30 and 4x = 2 - y then find x + y.

OpenStudy (anonymous):

6x-3y-30=0 4x+y-2=0 ----------- 6x-3x-30=0 12x+3y-6=0 + ------------ 18x=36 x=2 ------------ 12*2+3y-6=0

OpenStudy (anonymous):

3y=-18 y=-6

OpenStudy (anonymous):

x+y=? 2+(-6)=-4

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok im confused where did 6x-3x-30 come from? 12+3y-6

OpenStudy (anonymous):

6x - 3y = 30 => 6x-3x-30=0 4x = 2 - y => 4x+y-2=0

OpenStudy (anonymous):

4x+y-2=0 (multiplying by 3) makes 12x+3y-6=0

OpenStudy (anonymous):

ok i was just confused because you keep typing 3x up at the top instead of 3y but i get it now thanks

OpenStudy (anonymous):

sry..now i see.. it was supposed to be 6x-3y-30=0

OpenStudy (anonymous):

You re-arrange the equation and you'll have 6x-3y-30=0 4x+y-2=0 Solve using elimination method in simultaneous equation multiply the first line by 6 and the second line by 4 and we'll have 4 (6x-3y-30)=0 6(4x+y-2)=0 = 24x -12y-120=0 ----Equation 1 24x +6y -12=0 ---Equation 2 Subtract equation 2 from equation 1 -18y-108=0 -18y=108 y= - 108/18 y= -6 Substitute the value for y in 6x-3y-30=0 to find x 6x-3(-6)-30=0 6x+18-30=0 6x-12=0 6x=12 x=12/6 x=2 x+y = 2+(-6)=2-6= -4 Hope this helps